OpenStudy (spareb665):

The radioactive element carbon‐14 has a half‐life of 5750 years. The percentage of carbon‐14 present in the remains of organic matter can be used to determine the age of that organic matter. Archaeologists discovered that the linen wrapping from one of the Dead Sea Scrolls had lost 15.8% of its carbon‐14 at the time it was found. How old was the linen wrapping?

OpenStudy (anonymous):

We know that N(t)=N0*(1/2)^(t/5750) So you can solve (1-0.158)=(1/2)^(t/575) for t to find the age

OpenStudy (anonymous):

Sorry meant 5750 there

OpenStudy (anonymous):

1426.620 years.